Table of Contents


SITE 1178 (ENT-09A)

The vessel reached Site ENT-09A (1178) at 0730 hr on 8 July, and preparations for making up the BHA began. Considerable current was observed, so the vessel was offset upcurrent 200 m from the location to drop the beacon. The beacon was dropped at 0845 hr on 8 July, establishing Site 1178. Once the hydrophones and thrusters were lowered and the ship settled out over the site in DP mode, it was determined that the beacon had come to rest on the seafloor 180 m downcurrent of the location.

Hole 1178A
Because of the strong current, the ship had to be offset upcurrent from the location and allowed to drift back to enable the drill collars to be made up. A BHA consisting of a 9-7/8-in PDC bit, bit sub, seal bore drill collar, landing saver sub, modified top sub, modified head sub, nonmagnetic drill collar, seven 8-1/4-in drill collars, a tapered drill collar, and crossover sub was assembled.
A PDR measurement was taken that indicated a drilling depth of 1716.5 mbsl (1728.4 mbrf). The bit was lowered to 1710 mbrf where a bottom water sample was collected using the WSTP. The bit was then lowered to 1728 mbrf; core was taken but no sediment was recovered. The bit was lowered to 1725.6 mbsl (1737.5 mbrf), and another core was taken, with no sediment recovered. The bit was once again lowered 9.5 m to 1735.1 mbsl (1747 mbrf), and Hole 1178A was spudded at 1855 hr on 8 July. Core 1H recovered 2.9 m of sediment, establishing the seafloor depth as 1741.7 mbsl (1753.6 mbrf).
Cores 1H to 44X were taken from 0 to 410.8 mbsf (1753.6—2164.4 mbrf), recovering 340.19 m of core (83%). The penetration rate had decreased to <6 m/hr, when the decision was made to terminate XCB coring and switch to RCB coring. Hole 1178A was filled with 122 bbl of heavy mud, and then the pipe was tripped out of the hole. The bit cleared the rig floor at 1430 hr on 11 July, officially ending Hole 1178A. Because Hole 1178A was the last deployment of the APC/XCB BHA for Leg 190, it was disassembled.

Hole 1178B
The ship was offset 20 m to the east as an RCB BHA was being assembled. The BHA consisted of a new 9-7/8-in CC-3 roller cone bit, a bit sub, 8-1/4-in drill collar, modified top sub, modified head sub, seven 8-1/4-in drill collars, a tapered drill collar, six joints of 5-1/2-in drill pipe, and a crossover sub to 5-in drill pipe. The bit was tripped to the seafloor, and Hole 1178B was spudded at 2015 hr on 11 July.
Hole 1178B was first drilled to 395 mbsf (1753.6—2148.6 mbrf). RCB coring then commenced, and Cores 2R through 31R were taken from 395.0 to 679.2 mbsf (2148.6—2432.8 mbrf), recovering 165.96 m of core (58%).
During operations at Site 1178, the current generally averaged between 2.5 and 3.5 nmi/hr. However, on several occasions it was believed that the current was over 5 nmi/hr. The bridge speed log indicated 4.8 nmi/hr during those times, and 100 turns on both screws were required to maintain station. Severe vibrations in the drill string were observed. The vibrations were transmitted up to the top drive and associated traveling equipment. The electrical supervisor reported that the hydrophones were vibrating severely within their wells. The aft side of the upper guide horn was too hot to touch with a bare hand because of the drill string rubbing against it.
Even though the ship was able to maintain position during the peak current times, we thought that the safe operational limits had been reached for two reasons. First, the automatic station keeping (ASK) system was already operating at its maximum output available from the main screws. Any increase in the 10—12 nmi/hr head winds would have exceeded the ASK systems ability to maintain station. Second, and perhaps more importantly, the severe vibration in the drill string and traveling equipment was of great concern. Fatiguing of drill-pipe connections as well as possible damage to compensator rods and seals was a distinct possibility. There were also concerns about any loose objects that might fall out of the derrick. Had the current increased above the perceived 5 nmi/hr, operations would have had to be shut down.
At 1515 hr on 12 July, three thrusters (T-2, T-4, and T-5) automatically tripped off line. Fortunately the ASK system was able to compensate for the lost thrusters and maintain ship position. With a 3+ nmi/hr current running at the time, the decision was made to raise the bit from the current total depth of the hole (419.3 mbsf) to 32.7 mbsf until the ASK system problem could be resolved. The problem was traced to a faulty transformer in the forward thyrig room. This transformer supplied power to the thruster controller cards and was giving a faulty signal. The transformer was bypassed, and the ASK system performance immediately returned to normal. Operations were resumed after losing 4.5 hr.
Coring was stopped in Hole 1178B when the allotted coring time for Leg 190 expired at 0600 hr on 15 July. The hole was filled with heavy mud and the bit was tripped out. The bit cleared the seafloor at 0845 hr, and the positioning beacon was released and recovered at 1045 hr. The bit cleared the rig floor at 1250 hr on 15 July, officially ending Hole 1178B and operations at Site 1178.

Transit from Site 1178 to Yokohama
The ship began the transit to Yokohama at 1400 hr on 15 July. The first line was ashore at 1600 hr on 16 July at Daikoku Wharf, Yokohama, Japan, officially ending Leg 190.

Table of Contents